# Meta-Unlearning on Diffusion Models: Preventing Relearning Unlearned Concepts

### v1

@@ 5.2 Copyright and style removal
In this section, we use both ESD-u-1-based unlearning and meta-unlearning to remove copyrighted concepts and/or styles from images.
Copyright removal. Fig. 2 shows the removal of two copyrighted cartoon characters, SpongeBob and Snoopy. After unlearning, neither unlearned nor meta-unlearned models are able to produce the removed copyright concepts. However, after maliciously finetuned on unlearned concepts, the unlearned model becomes to relearn the unlearned concepts, while the meta-unlearned model is resistant to relearning and avoids copyright infringement.
[…]
Figure 5: The value of orthogonal term ${\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)}$ for each step of meta-unlearning. Because the value is noisy, we use the regression line to represent a smoothed trend.
@@
Performance on unrelated concepts. Fig. 9 displays the images generated on prompts unrelated to the removed concepts before the unlearned or meta-unlearned model is maliciously finetuned. This demonstrates that both unlearned and meta-unlearned models can maintain the capability to generate images on unrelated and benign concepts.
@@ 5.3 More analyses
In this section, we first discuss the relationship between the unlearn concept and its related concept during meta-unlearning and malicious finetuning. Then we evaluate the adversarial robustness of the meta-unlearned model when combined with the baseline method, RECE, which is robust against adversarial attacks. Refer to Appendix B for the performance of our method under more metrics.
@@
Robustness to adversarial attacks. Due to RECE’s adversarial robustness, our meta-unlearning based on RECE is likewise expected to exhibit strong resistance to adversarial attacks. We utilize UnlearnDiffAtk (Zhang et al., 2023) as the evaluation framework for adversarial robustness. Compared with the attack successful rate (ASR) for the RECE unlearned model, 64.79%, the ASR on our meta-unlearned model achieves 62.68%. Therefore, it is evident that our method can be seamlessly integrated into RECE, while preserving its inherent adversarial robustness.
@@ 6 Conclusion
In this paper, we present a meta-unlearning framework for DMs that effectively prevents the relearning of previously unlearned concepts, such as harmful content, copyright, and image style. Our method combines a meta objective with existing unlearning methods, ensuring that if a model is maliciously finetuned on unlearned data, related benign concepts self-destruct, impeding the relearning process. Extensive experiments on SD-v1-4 and SDXL reveal that our method maintains generation quality on benign data while significantly reducing the ability to generate unlearned concepts, even after adversarial finetuning. Our framework is compatible with a variety of unlearning techniques and offers a simple yet effective solution for improving the safety of DMs against potential misuse.
@@ Appendix A More related work
Meta-learning is generally used in few-shot learning to enhance performance by learning shared features from other data. The metric-based (Snell et al., 2017) and model-based meta-learning methods (Mishra et al., 2017; Munkhdalai & Yu, 2017; Santoro et al., 2016) rely on extra features or models to improve the few-shot learning capabilities. Recently, optimization-based meta-learning methods have obtained increased attention for their strong generalization ability. The optimization-based methods reduce the meta-learning problem into a bi-level optimization problem. The inner loop optimizes the base model on a certain task, and the outer loop optimizes the base model across several tasks to adjust the initial weight for quick adaption. Without introducing new elements, such a structure has the potential to adapt better to unseen data. MAML (Finn et al., 2017) is the most representative optimization-based method. Subsequent MAML variants (Rajeswaran et al., 2019; Nichol, 2018; Lee et al., 2019; Rusu et al., 2018) focus on optimizing the optimization process. Recent works (Henderson et al., 2023; Tamirisa et al., 2024) also proposed some meta-learning approaches for robustly preventing models from learning harmful tasks in language models.
@@ Appendix B Evaluation on more metrics
To further demonstrate the superiority of our method compared to the baseline, we conduct evaluation on ESD-f-3 unlearned and meta-unlearned SD-v-1-4 models with two metrics: Unsafe score and NSFW score. The Unsafe score is calculated as the percentage of images deemed harmful by SD’s safety checker (Rombach et al., 2022). The NSFW score is the average harmfulness score for each image, determined using Laion’s CLIP-based detector [^1] https://github.com/LAION-AI/CLIP-based-NSFW-Detector. We use the prompts of nudity subset in I2P dataset the, the same as those used in evaluation experiment in section 5. Although these two metrics assess general NSFW content rather than specifically targeting nudity, table 3 still illustrates that after malicious finetuning, the meta-unlearned SD exhibits a lower level of harmfulness compared to the unlearned SD.
@@ E.2 Hyperparameter
Following the papers of ESD (Gandikota et al., 2023) and SDD (Kim et al., 2023), we train ESD-based meta-unlearned model and SDD-based meta-unlearned model for 1000 and 1500 steps separately. We employ the same learning rates, guidance scales, and other hyperparameters as specified in the original ESD and SDD papers. The $\gamma_{2}$ in meta-unlearning is set to 0.05 for ESD-u-1, and to 0.1 for ESD-u-3, ESD-f-3, and SDD, respectively. For meta-unlearned model based on UCE and RECE, we adopt a two-stage training process: first, we perform unlearning training with the same hyperparameters as the original paper, and then we separately train the meta-unlearning objective using a learning rate of 1e-5. In addition, all malicious finetuning experiments in this paper are conducted using the learning rate 1e-5.

### v2

@@ 5.2 Copyright and style removal
We use ESD-u-1-based unlearning and meta-unlearning to remove copyright/style concepts from images.
Copyright removal. Fig. 2 shows removal of two copyrighted cartoon characters, SpongeBob and Snoopy. After unlearning, neither unlearned nor meta-unlearned models are able to produce the removed copyright concepts. However, after maliciously finetuned on unlearned concepts, the unlearned model becomes to relearn the unlearned concepts, while the meta-unlearned model is resistant to relearning and avoids copyright infringement.
@@
Performance on unrelated concepts. Fig. 9 (in Appendix) displays images generated on prompts unrelated to removed concepts before the unlearned or meta-unlearned model is maliciously finetuned. This shows that both unlearned and meta-unlearned models can maintain the capability to generate images on unrelated/benign concepts.
@@ 5.3 More analyses
In this section, we first discuss the relationship between the unlearn concept and its related concept during meta-unlearning and malicious finetuning. Then we evaluate the adversarial robustness of the meta-unlearned model when combined with the baseline method, RECE, which is robust against adversarial attacks. Refer to Appendix B for the performance of our method under more metrics and Appendix F for hyperparameter analysis.
@@
Robustness to adversarial attacks. Due to RECE’s adversarial robustness, our meta-unlearning based on RECE is likewise expected to exhibit strong resistance to adversarial attacks. We utilize UnlearnDiffAtk [73] as the evaluation framework for adversarial robustness. Compared with the attack successful rate (ASR) for the RECE unlearned model, 64.79%, the ASR on our meta-unlearned model achieves 62.68%. Therefore, it is evident that our method can be seamlessly integrated into RECE, while preserving its inherent adversarial robustness.
@@ 6 Conclusion
Our meta-unlearning framework combines a meta objective with existing unlearning methods, ensuring that if an already unlearned model is maliciously finetuned on unlearned data, related benign concepts will self-destruct, impeding the relearning process. Extensive experiments on SD-v1-4 and SDXL reveal that our method maintains benign generation quality while largely reducing generation of unlearned concepts. Meta-unlearning is compatible with a variety of unlearning techniques and offers a simple yet effective solution for preventing DMs from potential misuse.
Acknowledgements
This work is supported in part by NSF of China (Nos. 92470118, 62306176), Natural Science Foundation of Shanghai (No. 23ZR1428700), CCF-ALIMAMA TECH Kangaroo Fund (NO. CCF-ALIMAMA OF 2025010), and CCF-Zhipu Large Model Innovation Fund (No. CCF-Zhipu202412).
@@ Appendix A More related work
A.1 Machine unlearning on language models
While this paper primarily focuses on unlearning DMs, there have been a lot of efforts devoted to unlearning language models [69, 35, 65, 28, 68, 11, 72, 22, 62, 60, 59]. These methods typically finetune the model on a forget set. In addition, there are also other tuning-free unlearning techniques, including contrastive decoding [19, 64, 21, 3], task vectors [4, 32], in-context learning [43, 38, 61], and input processing and detection [1, 8, 31].
A.2 Meta-learning
Meta-learning is generally used in few-shot learning to enhance performance by learning shared features from other data. The metric-based [56] and model-based meta-learning methods [36, 37, 53] rely on extra features or models to improve the few-shot learning capabilities. Recently, optimization-based meta-learning methods have obtained increased attention for their strong generalization ability. The optimization-based methods reduce the meta-learning problem into a bi-level optimization problem. The inner loop optimizes the base model on a certain task, and the outer loop optimizes the base model across several tasks to adjust the initial weight for quick adaption. Without introducing new elements, such a structure has the potential to adapt better to unseen data. MAML [5] is the most representative optimization-based method. Subsequent MAML variants [49, 39, 27, 52] focus on optimizing the optimization process. Recent works [13, 59] also proposed some meta-learning approaches for robustly preventing models from learning harmful tasks in language models.
@@ Appendix B Evaluation on more metrics
To further demonstrate the superiority of our method compared to the baseline, we conduct evaluation on ESD-f-3 unlearned and meta-unlearned SD-v-1-4 models with two metrics: Unsafe score and NSFW score. The Unsafe score is calculated as the percentage of images deemed harmful by SD’s safety checker [51]. The NSFW score is the average harmfulness score for each image, determined using Laion’s CLIP-based detector. [^1] https://github.com/LAION-AI/CLIP-based-NSFW-Detector We use the prompts of nudity subset in I2P dataset the, the same as those used in evaluation experiment in Section 5. Although these two metrics assess general NSFW content rather than specifically targeting nudity, table 3 still illustrates that after malicious finetuning, the meta-unlearned SD exhibits a lower level of harmfulness compared to the unlearned SD.
@@ E.2 Hyperparameter
Following the papers of ESD [6] and SDD [24], we train ESD-based meta-unlearned model and SDD-based meta-unlearned model for 1000 and 1500 steps separately. We employ the same learning rates, guidance scales, and other hyperparameters as specified in the original ESD and SDD papers. The $\gamma_{2}$ in meta-unlearning is set to 0.05 for ESD-u-1, and to 0.1 for ESD-u-3, ESD-f-3, and SDD, respectively. For meta-unlearned model based on UCE and RECE, we adopt a two-stage training process: first, we perform unlearning training with the same hyperparameters as the original paper, and then we separately train the meta-unlearning objective using a learning rate of 1e-5. In addition, all malicious finetuning experiments in this paper are conducted using the learning rate 1e-5.
Appendix F Analysis of Hyperparameters
Hyperparameters of $\gamma_{1}$ and $\gamma_{2}$.
When training is insufficiently saturated, increasing $\gamma_{1}$ improves the removal of unsafe content before malicious finetuning (FT) but weakens resistance to it. Increasing $\gamma_{2}$ enhances safety after malicious FT but reduces the effect of initial unsafe content removal. With sufficient training steps, the $\gamma_{1}$ to $\gamma_{2}$ ratio becomes less significant, ultimately achieving the same effect. Table 4 shows the results of varying $\gamma_{1}$ and $\gamma_{2}$ ratios.
Table 4: Nudity score of various $\gamma_{1}$ and $\gamma_{2}$ ratios for ESD-u-1.
[…]
More commonly used update rules.
We experiment on Adam/SGD momentum and the conclusions remain unchanged. Taking ESD-u-3 as an example, the nudity scores after malicious FT were 26.76, 25.35, and 26.06 for SGD, Adam, and SGD momentum. More results will be included.
Different values of $M$.
As seen in Table 5, larger $M$ makes the model better generate harmless content before malicious FT but weakens its resistance to malicious FT.
Table 5: Transposed results of varying $M$ for ESD-u-3.
M 1 3 5 FID (Before FT) $\downarrow$ 20.52 19.72 17.92 CLIPScore (Before FT) $\uparrow$ 29.65 30.36 30.98 Nudity Score (After FT) $\downarrow$ 26.76 28.17 32.39